# Approximate independence of permutation mixtures

### v2

@@ 1 Introduction
$\left(\mathbb{E}_{\mathbb{P}_{n}}[f]-\mathbb{E}_{\mathbb{Q}_{n}}[f]\right)^{2}\leq\mathbb{E}_{\overline{P}^{\otimes n}}\left[f^{2}\right]\cdot\frac{12{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{3{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}}{n}.$
Theorem 1.6 shows that the squared TV distance between $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ is of the order $O_{\mathcal{P}}(\frac{1}{n})$. In addition, the mean difference of any test function $f$ under $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ is of order $O_{\mathcal{P}}(n^{-1/2})$ as long as the second moment of $f$ under the i.i.d. distribution $\overline{P}^{\otimes n}$ is bounded.
This result turns out to have implications in empirical Bayes models. In the “oracle Bayes” setting of empirical Bayes [GR09, Efr19], a distribution class $(P_{\theta})_{\theta\in\Theta}$ and a set of parameters $\theta:=\{\theta_{1},\cdots,\theta_{n}\}$ are given. The true parameter $\vartheta$ is a uniformly random permutation of $\theta$, i.e. $\vartheta_{i}=\theta_{\pi(i)}$ for all $i\in[n]$ with $\pi\sim\text{\rm Unif}(S_{n})$, and the observation $X=(X_{1},\dots,X_{n})$ follows the distribution $\otimes_{i=1}^{n}P_{\vartheta_{i}}$ conditioned on $\vartheta$. To estimate $\vartheta$, there are two natural types of decision rules to compete with, following the language of [GR09]: one is a simple decision rule, taking the form $\widehat{\vartheta}^{\mathrm{S}}=(\Delta(X_{1}),\dots,\Delta(X_{n}))$ for some fixed function $\Delta$, and the other is a permutation invariant decision rule, satisfying
$\widehat{\vartheta}^{\mathrm{PI}}_{\pi(i)}(X_{\pi(1)},\dots,X_{\pi(n)})=\widehat{\vartheta}^{\mathrm{PI}}_{i}(X_{1},\dots,X_{n}),\quad i\in[n],$
for all $\pi\in S_{n}$. Denote by ${\mathcal{D}}^{\mathrm{S}}$ and ${\mathcal{D}}^{\mathrm{PI}}$ the classes of all simple and permutation invariant decision rules, respectively. Given a separable loss function $L(\vartheta,\widehat{\vartheta})=\sum_{i=1}^{n}\ell(\vartheta_{i},\widehat{\vartheta}_{i})$, we define the optimal risks among all decision rules in ${\mathcal{D}}^{\mathrm{S}}$ and ${\mathcal{D}}^{\mathrm{PI}}$:
$r^{\mathrm{S}}(\theta)=\inf_{\widehat{\vartheta}\in{\mathcal{D}}^{\mathrm{S}}}\mathbb{E}\left[L(\vartheta,\widehat{\vartheta})\right],\qquad r^{\mathrm{PI}}(\theta)=\inf_{\widehat{\vartheta}\in{\mathcal{D}}^{\mathrm{PI}}}\mathbb{E}\left[L(\vartheta,\widehat{\vartheta})\right].$
As ${\mathcal{D}}^{\mathrm{S}}\subseteq{\mathcal{D}}^{\mathrm{PI}}$, we always have $r^{\mathrm{S}}(\theta)\geq r^{\mathrm{PI}}(\theta)$. Theorem 1.6 implies that the risk difference $r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)$ is small (i.e. of the order $o(n)$) in general:
Corollary 1.7.
$r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)\leq M\sqrt{6n{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{3{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}}.$
Corollary 1.7 shows that for a large class of models and losses, asymptotically these two types of decision rules achieve similar mean risks $r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)=O_{{\mathcal{P}}}(\sqrt{n})$, strengthening a “folklore” result $r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)=o(n)$ in [HR55]. For the quadratic loss $L(\vartheta,\widehat{\vartheta})=\|\vartheta-\widehat{\vartheta}\|_{2}^{2}$, the risk difference can be improved to $O(1)$ via a different strategy, as shown for a class of models in [GR09]; we will discuss this improvement in Section 5.3.
@@ 1.1 Notation
Throughout the paper all logarithms are in base $e$. For a positive integer $n$, let $[n]:=\{1,\ldots,n\}$, and $S_{n}$ be the permutation group over $[n]$. For a vector $x$ in $\mathbb{R}^{n}$ or $\mathbb{C}^{n}$, let $x^{\top}$ and $x^{\mathsf{H}}$ be its real and conjugate transposes, respectively, and $e_{\ell}(x):=\sum_{S\subseteq[n]:|S|=\ell}\prod_{i\in S}x_{i}$ be the elementary symmetric polynomial of degree $\ell$ in $x$. For a square matrix $A=(a_{ij})_{i,j\in[n]}\in\mathbb{R}^{n\times n}$, let $\mathrm{Tr}(A)=\sum_{i=1}^{n}a_{ii}$ be its trace, and $\mathrm{Perm}(A)=\sum_{\pi\in S_{n}}\prod_{i=1}^{n}a_{i\pi(i)}$ be its permanent. Let ${\mathcal{N}}(\mu,\sigma^{2})$ denote the normal distribution with mean $\mu$ and variance $\sigma^{2}$, and ${\mathcal{C}}{\mathcal{N}}(0,1)$ denote the complex normal distribution with real and imaginary parts being independent ${\mathcal{N}}(0,\frac{1}{2})$ random variables. For probability measures $\mu$ and $\nu$, let $L^{2}(\mu)$ be the space of all functions $f$ with $\mathbb{E}_{\mu}\left[f^{2}\right]<\infty$, $\mu\otimes\nu$ be the product measure, and $\mu\star\nu$ be the convolution defined as $\mu\star\nu(A)=\int\mu(A-x)\mathrm{d}\nu(x)$. For probability measures $P$ and $Q$ on the same probability space, let
${\sf TV}(P,Q)=\frac{1}{2}\int|\mathrm{d}P-\mathrm{d}Q|,\quad H^{2}(P,Q)=\int\left(\sqrt{\mathrm{d}P}-\sqrt{\mathrm{d}Q}\right)^{2},\quad\mathsf{LC}(P,Q)=\frac{1}{2}\int\frac{(\mathrm{d}P-\mathrm{d}Q)^{2}}{\mathrm{d}P+\mathrm{d}Q}$
be the total variation (TV), squared Hellinger, and Le Cam distances, respectively, and
@@
be the Kullback–Leibler (KL) and $\chi^{2}$ divergences, respectively. A collection of inequalities between the above distances/divergences can be found in [PW24, Chapter 7.6].
@@ 1.2 Proof techniques
Both applications in Theorem 1.5 and 1.6 rely on a combination of the above two approaches, by first expanding $M(x)$ around the all-ones matrix $J$ and then relating each term in the expansion to a counterpart appearing in the expression of $\mathrm{Perm}(A)$, through the identities in Appendix A. In particular, a complex version of the inequality (10) for $x\in\mathbb{C}^{n}$ turns out to be crucial in the proof of Theorem 1.6, a technique which could be of independent interest.
@@ 1.3 Related work
Comparisons between $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ first implicitly arose in the consideration of compound decision problems. An early result in this direction is due to Hannan [Han53] (see [HR55]). In our notation, his results reads as follows: suppose that ${\mathcal{P}}=\{P_{-},P_{+}\}$ consists of two elements, which are not orthogonal to each other. Given natural numbers $n$ and $k$ with $k\leq n$, denote by $\mathbb{P}^{(k)}_{n}$ the permutation mixture arising from taking $P_{1}=\dots=P_{k}=P_{+}$ and $P_{k+1}=\dots=P_{n}=P_{-}$. If $k/n$ is bounded away from 00 and $1$ as $n,k\to\infty$, then
@@
In [HR55], this result is used to establish asymptotic equivalence between the performance of permutation invariant and simple decision rules. Subsequent empirical Bayes literature largely focused on simple decision rules and avoided explicit consideration of $\mathbb{P}_{n}$, but the importance of comparing $\mathbb{P}_{n}$ to $\mathbb{Q}_{n}$ to obtain genuine oracle inequalities was emphasized by Greenshtein and Ritov [GR09, GR19], who obtained risk bounds for the two models under the square loss. The general power of permutation-invariant decision rules was recently investigated in [Wei21].
The fact that exchangeable distributions are “essentially” (mixtures of) product measures is an important heuristic which has its roots in the work of de Finetti [DF29] and Hewitt–Savage [HS55]. Diaconis and Friedman observed that versions of this claim hold in a number of different models [DF87], and precise quantitative versions have been established for finite exchangeable sequences under different assumptions [Sta78, DF80, GK21, JGK24]. Recently, motivated by applications in conformal prediction (see, e.g., [TBCR19, BCRT23a]), similar results have been obtained for “weighted” exchangeable sequences as well [BCRT23b].
Bounding the $\chi^{2}$ divergence for mixture distributions is a common task in high-dimensional statistics [LNS99, Bar02, CL11, IS12, WY16, CCT17, BMV+18, JHW18, PWBM18, BW19, BNWR20], and obtaining bounds for “low-degree” versions of the $\chi^{2}$ divergence has become an important technique in establishing statistical-computational gaps (see, e.g., [KWB19]). In most of these applications, $\mathbb{P}_{n}$ is a mixture distribution but $\mathbb{Q}_{n}$ is particularly simple (for instance, its coordinates are i.i.d. ${\mathcal{N}}(0,1)$ or $\text{\rm Bern}(1/2)$); in this case, expanding the likelihood ratio $\frac{\mathrm{d}\mathbb{P}_{n}}{\mathrm{d}\mathbb{Q}_{n}}$ in a basis of $L^{2}(\mathbb{Q}_{n})$-orthogonal polynomials yields explicit bounds in terms of the moments of $\mathbb{P}_{n}$. Few general techniques exist to obtain sharp results outside of this simple setting, though Schramm and Wein [SW22] (see also [RSWY23]) showed that a related expansion based on cumulants yields a (potentially loose) bound when $\mathbb{Q}_{n}$ is a mixture distribution. In a more general context, Kunisky [Kun24] has shown that employing the Hoeffding decomposition [Hoe48] can sometimes obtain comparable results for more complicated choices of $\mathbb{Q}_{n}$. We employ a similar technique in Section 3.
@@
He also developed analogous bounds for a Bernoulli version of the problem. A dimension-independent upper bound on the KL divergence ${\sf KL}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ was also observed and obtained in [TP23, Theorem 2] for a finite class ${\mathcal{P}}$, used as a crucial technical step towards establishing the capacity upper bound of the noisy permutation channel in [Mak20]. Their proof relies on an anti-concentration property for the histogram of independent discrete observations with non-identical distributions, leading to an upper bound linear in $|{\mathcal{P}}|$. They also prove a bound analogous to the one presented in Theorem 1.5 for $k$-dimensional marginals with $k<n$ via a convexity argument [TP23, Appendix A]; however, their bound scales as $O(\frac{k}{n})$ rather than $O(\frac{k^{2}}{n^{2}})$.

### v3

@@ 1 Introduction
$\left(\mathbb{E}_{\mathbb{P}_{n}}[f]-\mathbb{E}_{\mathbb{P}^{\prime}_{n}}[f]\right)^{2}\leq\mathbb{E}_{\overline{P}^{\otimes n}}\left[f^{2}\right]\cdot\frac{12{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{3{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}}{n}.$
Theorem˜1.5 shows that the squared TV distance between $\mathbb{P}_{n}$ and $\mathbb{P}^{\prime}_{n}$ is of the order $O_{\mathcal{P}}(\frac{1}{n})$. In addition, the mean difference of any test function $f$ under $\mathbb{P}_{n}$ and $\mathbb{P}^{\prime}_{n}$ is of order $O_{\mathcal{P}}(n^{-1/2})$ as long as the second moment of $f$ under the i.i.d. distribution $\overline{P}^{\otimes n}$ is bounded. This result will find applications in compound decision problems and differential privacy in Section˜2.
@@ 1.1 Notation
Throughout the paper all logarithms are in base $e$. For a positive integer $n$, let $n]:=\{1,\ldots,n\}$, and $S_{n}$ be the permutation group over $[n]$. For a vector $x$ in $\mathbb{R}^{n}$ or $\mathbb{C}^{n}$, let $x^{\top}$ and $x^{\mathsf{H}}$ be its real and conjugate transposes, respectively, and $e_{\ell}(x):=\sum_{S\subseteq[n]:|S|=\ell}\prod_{i\in S}x_{i}$ be the elementary symmetric polynomial of degree $\ell$ in $x$. For a square matrix $A=(a_{ij})_{i,j\in[n]}\in\mathbb{R}^{n\times n}$, let $\mathrm{Tr}(A)=\sum_{i=1}^{n}a_{ii}$ be its trace, and $\mathrm{Perm}(A)=\sum_{\pi\in S_{n}}\prod_{i=1}^{n}a_{i\pi(i)}$ be its permanent. Let ${\mathcal{N}}(\mu,\sigma^{2})$ denote the normal distribution with mean $\mu$ and variance $\sigma^{2}$, and ${\mathcal{C}}{\mathcal{N}}(0,1)$ denote the complex normal distribution with real and imaginary parts being independent ${\mathcal{N}}(0,\frac{1}{2})$ random variables. For probability measures $\mu$ and $\nu$, let $L^{2}(\mu)$ be the space of all functions $f$ with $\mathbb{E}_{\mu}\left[f^{2}\right]<\infty$, $\mu\otimes\nu$ be the product measure, and $\mu\star\nu$ be the convolution defined as $\mu\star\nu(A)=\int\mu(A-x)\mathrm{d}\nu(x)$. For probability measures $P$ and $Q$ on the same probability space, let
${\sf TV}(P,Q)=\frac{1}{2}\int|\mathrm{d}P-\mathrm{d}Q|,\qquad H^{2}(P,Q)=\int\left(\sqrt{\mathrm{d}P}-\sqrt{\mathrm{d}Q}\right)^{2}$
be the total variation (TV) and squared Hellinger distances, respectively, and
@@
be the Kullback–Leibler (KL) and $\chi^{2}$ divergences, respectively. A collection of inequalities between the above distances/divergences can be found in [PW25, Chapter 7.6].
@@ 1.2 Proof techniques
Both applications in Theorem˜1.4 and 1.5 rely on a combination of the above two approaches, by first expanding $M(x)$ around the all-ones matrix $J$ and then relating each term in the expansion to a counterpart appearing in the expression of $\mathrm{Perm}(A)$, through the identities in Appendix˜A. In particular, a complex version of the inequality (8) for $x\in\mathbb{C}^{n}$ turns out to be crucial in the proof of Theorem˜1.5, a technique which could be of independent interest.
@@ 1.3 Related work
Comparisons between $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ first implicitly arose in the consideration of compound decision problems. An early result in this direction is due to Hannan [Han53] (see [HR55]). In our notation, his results reads as follows: suppose that ${\mathcal{P}}=\{P_{-},P_{+}\}$ consists of two elements, which are not orthogonal to each other. Given natural numbers $n$ and $k$ with $k\leq n$, denote by $\mathbb{P}^{(k)}_{n}$ the permutation mixture arising from taking $P_{1}=\dots=P_{k}=P_{+}$ and $P_{k+1}=\dots=P_{n}=P_{-}$. If $k/n$ is bounded away from 0 and $1$ as $n,k\to\infty$, then
@@
In [HR55], this result is used to establish asymptotic equivalence between the performance of permutation invariant and simple decision rules. Subsequent empirical Bayes literature largely focused on simple decision rules and avoided explicit consideration of $\mathbb{P}_{n}$, but the importance of comparing $\mathbb{P}_{n}$ to $\mathbb{Q}_{n}$ to obtain genuine oracle inequalities was emphasized by Greenshtein and Ritov [GR09, GR19], who obtained risk bounds for the two models under the squared loss. The general power of permutation-invariant decision rules was recently investigated in [Wei21].
The fact that exchangeable distributions are “essentially” (mixtures of) product measures is an important heuristic which has its roots in the work of de Finetti [DF29] and Hewitt–Savage [HS55]. Diaconis and Friedman observed that versions of this claim hold in a number of different models [DF87], and precise quantitative versions have been established for finite exchangeable sequences under different assumptions [Sta78, DF80, Bob05, Roo15, GK21, JGI25]. Recently, motivated by applications in conformal prediction (see, e.g., [TBCR19, BCRT23]), similar results have been obtained for “weighted” exchangeable sequences as well [BCRT24].
Bounding the $\chi^{2}$ divergence for mixture distributions is a common task in high-dimensional statistics [LNS99, Bar02, CL11, IS12, WY16, CCT17, BMV+18, JHW18, PWBM18, BW19, BNWR20], and obtaining bounds for “low-degree” versions of the $\chi^{2}$ divergence has become an important technique in establishing statistical-computational gaps (see, e.g., [KWB19]). In most of these applications, $\mathbb{P}_{n}$ is a mixture distribution but $\mathbb{Q}_{n}$ is particularly simple (for instance, its coordinates are i.i.d. ${\mathcal{N}}(0,1)$ or $\text{\rm Bern}(1/2)$); in this case, expanding the likelihood ratio $\frac{\mathrm{d}\mathbb{P}_{n}}{\mathrm{d}\mathbb{Q}_{n}}$ in a basis of $L^{2}(\mathbb{Q}_{n})$-orthogonal polynomials yields explicit bounds in terms of the moments of $\mathbb{P}_{n}$. Few general techniques exist to obtain sharp results outside of this simple setting, though Schramm and Wein [SW22] (see also [RSWY23]) showed that a related expansion based on cumulants yields a (potentially loose) bound when $\mathbb{Q}_{n}$ is a mixture distribution. In a more general context, Kunisky [Kun25] has shown that employing the Hoeffding decomposition [Hoe48] can sometimes obtain comparable results for more complicated choices of $\mathbb{Q}_{n}$. We employ a similar technique in Section˜4.
@@
He also developed analogous bounds for a Bernoulli version of the problem. A dimension-independent upper bound on the KL divergence ${\sf KL}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ was also observed and obtained in [TP23, Theorem 2] for a finite class $\mathcal{P}$}, used as a crucial technical step towards establishing the capacity upper bound of the noisy permutation channel in [Mak20]. Their proof relies on an anti-concentration property for the histogram of independent discrete observations with non-identical distributions, leading to an upper bound linear in $|{\mathcal{P}}|$. They also prove a bound analogous to the one presented in Theorem˜1.4 for $k$-dimensional marginals with $k<n$ via a convexity argument [TP23, Appendix A]; however, their bound scales as $O(\frac{k}{n})$ rather than $O(\frac{k^{2}}{n^{2}})$.